Twenty years ago the age of the father was four times the age of son . After four years from now the age of the father will be double that of son .Give algebraic and graphical representation of the situation and hence find their ages.

Answer by mananth(16946)

Twenty years ago the age of the father was four times the age of son . After four years from now the age of the father will be double that of son .Give algebraic and graphical representation of the situation and hence find their ages.

let sons age now be x
20 years ago son's age = x-20
fathers age = 4(x-20)
4 years hence
sons age = x+4
fathers age = 4(x-20)+24
4(x-20)+24= 2(x+4)
4x-80+24=2x+8
4x-56=2x+8
4x-2x=56+8
2x=64
x=32 son's age now
20 years ago father's age = 4(x-20)=>4(32-20)=>48
So now father is 68 years